OpenStudy (anonymous):

Find an equation of the tangent line to y=(x^2)-3 at x=2.

terenzreignz (terenzreignz):

Tangent lines... you mind should instantly jump to 'derivatives' :D to get any line at all, you need its slope, and a point on that line. The slope, you can find, by finding the derivative of this function... which is...?

OpenStudy (anonymous):

yup i'm doing derivatives in calc one right now. so to find the derivative I should plug in the two?

OpenStudy (whpalmer4):

No, you find the derivative function with the power rule, then evaluate it at x = 2 to get the slope of the tangent line at x=2...

terenzreignz (terenzreignz):

First find the derivative.

OpenStudy (whpalmer4):

Then you'll fit a line through the point (2,y(2)) with the slope you found.

OpenStudy (whpalmer4):

Do you have the derivative yet?

OpenStudy (whpalmer4):

Happy to help, but you need to do something...

OpenStudy (anonymous):

sorry I was doing something else but i'm back. ok so I'll try to find the derivative first

OpenStudy (whpalmer4):

Good!

OpenStudy (anonymous):

ok so not quite sure if I did this right in finding the derivative

OpenStudy (anonymous):

\[x ^{2}+2x \Delta x-3+\Delta x ^{2}-3\div \Delta x\]

OpenStudy (anonymous):

that was by using the equation = f(x+Δx) - f(x) / deltax

OpenStudy (anonymous):

i know I need to simplify it further... to 2xdeltax-6+deltax. but I thinking something is wrong there.

OpenStudy (whpalmer4):

aren't you taking a limit?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

cant you just use the power rule? \[nx^{n-1}\]

OpenStudy (anonymous):

or in your case \[ax^n =anx^{n-1}\]

OpenStudy (whpalmer4):

what I was referring to is the definition of the derivative as \[y'(x) = \lim_{\Delta x->0}\frac{f(x+\Delta x)-f(x)}{\Delta x}\]

OpenStudy (anonymous):

ah i see will if limits are required by teacher to solving this answer ignore me :)

OpenStudy (whpalmer4):

I would use the power rule, if you know it, but if you don't...

OpenStudy (whpalmer4):

I doubt there's any requirement to do it the hard way, just that they may not have learned the easy way yet...

OpenStudy (anonymous):

ah i see well either way you get y prime so do what ever floats your boat :P

OpenStudy (whpalmer4):

but here, we can do it the hard way, it isn't all that hard for this function! Before we do, can you confirm that the equation is \[y = x^2-3\] and not \[y = (x-3)^2\]

OpenStudy (anonymous):

yes its the first one.

OpenStudy (anonymous):

yup

OpenStudy (whpalmer4):

\[y'(x) = \lim_{\Delta x->0}\frac{f(x+\Delta x) - f(x)}{\Delta x}\]\[y'(x) = \lim_{\Delta x->0}\frac{(x+\Delta x)^2-3 - (x^2-3)}{\Delta x}\]\[y'(x) = \lim_{\Delta x->0}\frac{(x^2+2x\Delta x + (\Delta x)^2-3 - (x^2-3)}{\Delta x}\]\[y'(x) = \lim_{\Delta x->0}\frac{(\cancel{x^2}+2x\Delta x + (\Delta x)^2\cancel{-3}\cancel{ - (x^2-3)}}{\Delta x} \]\[=\lim_{\Delta x->0} \frac{2x \Delta x + (\Delta x)^2}{\Delta x} = \lim_{\Delta x->0}(2x + \Delta x) = 2x\]

OpenStudy (whpalmer4):

Okay, so the slope of the tangent to our curve at any point along the curve can be found by evaluating \(y'(x) = 2x\). We want to do this at \(x = 2\). \(y'(2) = 2(2) = 4\) So our tangent line has a slope of 4. Now we use the point-slope formula for a line with known slope \(m\) passing through a point \((x_0,y_0))\) to write the equation for the tangent line at \(x=2\). \[y-y_0 = m(x-x_0)\] At \(x=2\), \(y = x^2 -3 = (2)^2-3 = 4-3 = 1\) Our known point is \((2,1)\) \[y - 1 = 4(x-2)\]\[y = 4x-8+1\]\[y = 4x-7\]is the equation of our tangent line at \(x = 2\). For your patience, a picture!

OpenStudy (anonymous):

wow your amazing! I was actually trying to work on it myself and that's the same thing I got! Thanks so much!

OpenStudy (whpalmer4):

here's a question: what is the slope of the tangent line at x = 0?